OpenStudy (anonymous):

An investment offers $15,000 per year for 25 years. If an investor can earn 12 percent annually on other investments, what is the current value of this investment? Round your answer to the nearest dollar.

OpenStudy (anonymous):

This is a question on Present Value of Annuity Factor. So the Present Value comes at 15000 x 7.843 = $ 117645.

OpenStudy (anonymous):

thanks to KallolSarkar698 \[pv = A \times( (1 - (1 \div (1+ r ) ^n)) / r)\] pv = present value of investment A= annual payment r = annual return rate n = years pv = 15000 x ((1 - (1 / 17.001) / 0.12) pv = 15000 x 7.843 = 117,647 $